# GenBFA: An Evolutionary Optimization Approach to Bit-Flip Attacks on LLMs

### v2

@@ Abstract
Current BFA techniques fail to exploit this vulnerability due to the difficulty of traversing the large parameter space to identify critical parameters. To address this, we propose AttentionBreaker, a novel framework specifically designed for LLMs that facilitates efficient parameter-space traversal to identify critical parameters. Additionally, we introduce GenBFA, an evolutionary optimization strategy, to further narrow down the critical parameter set and identify the most vulnerable bits, facilitating a more efficient attack. Empirical findings indicate the pronounced vulnerability of LLMs to AttentionBreaker: for instance, perturbing merely $3$ bits ($4.129\times 10^{-9}\%$ of total parameters) in LLaMA3-8B 8-bit weight quantized (W8) model results in total performance collapse, with accuracy on Massive Multitask Language Understanding (MMLU) tasks plummeting from $67.3\%$ to $0\%$ and perplexity escalating from $12.6$ to $4.72\times 10^{5}$. These results highlight the framework’s efficacy in exploiting and exposing inherent vulnerabilities in LLM architectures.
@@ 6.1 Experimental Setup
(a) LLaMA3-8B-Instruct W8.
(b) LLaMA3-8B-Instruct W4.
(c) BitNet W1.58.
Figure 6: Comparison of BFA on bits selected at random with the ones selected using AttentionBreaker on LLaMA3-8B-Instruct W8, W4, and W1.58 models. Metrics are calculated using the MMLU benchmark.
@@ 6.2.1 AttentionBreaker on 8-bit Quantized LLMs
Here, we present the efficacy of the proposed AttentionBreaker framework on 8-bit weight quantized (W8) LLMs. Table II compares the effects of Random Bit-Flip Attack (Random BFA) and AttentionBreaker on 8-bit versions of LLM across several benchmarks. For each model, the precision, number of parameters, and bit-flip Count required to disrupt performance are listed in the Table. Benchmark results show that AttentionBreaker consistently reduces accuracy to zero across datasets, proving more efficient than Random BFA in degrading model performance with fewer bit-flips. The results highlight the impact of bit-flips on key model performance metrics, including perplexity, MMLU, and LM harness benchmarks.
For instance, LLaMA3-8B-Instruct W8 suffers accuracy drop to 00% on MMLU tasks from the original accuracy of $67.3$% by merely 3 bit-flips ($\mathbf{4.129\times 10^{-9}}$bit-flip rate) as captured in Figure 6(a). At the same time, the perplexity measure increases from $12.14$ to $4.72\times 10^{5}$ post-attack. In comparison, random bit-flips require $\textbf{10}^{\textbf{7}}\times$ ($\approx\textbf{0.1}$% bit-flip rate) more bit-flips to accomplish a similar level of model degradation. A similar trend is observed for Phi-3-mini-128k-Instruct W8 models as well.
@@ 6.2.2 AttentionBreaker on 4-bit quantized LLMs
In this section, we present the performance of AttentionBreaker on 4-bit weight quantized (W4) LLMs. As shown in Figure 6(b), the attack on LLaMA3-8B W4 demonstrates that $\mathbf{7.8125\times 10^{-8}}$or 28 bit-flips results in a perplexity increase from $12.60$ to $3.8$×$\times 10^{4}$, with the MMLU score dropping from $60.3$% to 00. A similar trend is observed for Phi-3-mini-128k-Instruct W4 models as well. This showcases the efficiency of the proposed AttentionBreaker framework.
@@ 6.2.3 AttentionBreaker on low-precision LLM
We further evaluate the effectiveness of our approach for ultra-low precision models, such as BitNet W1.58[26]. BitNet W1.58 represents a 1.58-bit variant of LLMs where each model parameter (or weight) is constrained to ternary values $\{-1,0,1\}$. Given their low precision, these models are intuitively expected to be inherently resistant, as a single bit-flip induced error has limited impact. Employing our approach, we observe that 45811 or $45811\text{ or }5\times 10^{-4}$bit-flip rate is required to degrade the model performance to near 00% effectively as demonstrated in Figure 6(c). In comparison, random bit-flips require $10^{5}\times$ ($\approx 1$% bit-flip rate) more bit-flips to accomplish a similar level of model degradation. This demonstrates the efficacy of AttentionBreaker in efficiently attacking low-precision models as well.
@@
Model Model type Precision Number of Parameters Random BFA (# of bit-flips) AttentionBreaker (# of bit-flips) WikiText-2 MMLU LM Harness VQAv2 TextVQA
@@ 6.2.4 AttentionBreaker on VLM
We demonstrate the efficacy of AttentionBreaker on multimodal models such as VLMs, which integrate a vision encoder with a language model to enable multimodal tasks, including image comprehension, alongside text-based tasks. When applying our attack methodology to the LLaVA1.6-7B W8 model, which comprises 7.57 billion parameters, we observe that a minimal bit-flip rate of $\mathbf{2.394\times 10^{-8}}$%—equivalent to only 15 bit-flips—causes a substantial degradation in performance. Specifically, the VQAv2 and TextVQA scores, initially at $81.4\%$ and $64.2\%$, respectively, drop to $0\%$ following the attack. In comparison, random bit-flips require ($\mathbf{10^{7}\approx 0.12}$bit-flip rate) more bit-flips to accomplish a similar level of model degradation. This result underscores the efficacy of the AttentionBreaker framework in exploiting latent vulnerabilities within multimodal LLMs as well.
@@ 6.4.1 Gradient-free AttentionBreaker
In this experiment, we investigate the effectiveness of AttentionBreaker in a gradient-free context, where gradient information is inaccessible at all stages. Here, the importance scores are computed solely based on weight magnitudes by setting $\alpha=0.0$. Our results demonstrate that this magnitude-based approach maintains the attack’s efficacy, as it accurately identifies critical weights, enabling a highly impactful attack. We observe that, without gradient information, setting $\alpha=0$ allows the model to be attacked with only $\mathbf{4.129\times 10^{-9}}\%$ bit-flips, reducing model accuracy to zero. This degradation matches the attack’s effectiveness at $\alpha=0.5$, as depicted in Figure 6(a). Subsequent analysis of the attack vector confirms that identical weights and bits are selected for attack across both settings, underscoring AttentionBreaker’s versatility and adaptability. This robustness is particularly significant as many defenses against BFAs assume that gradient access is essential for an attacker, prompting defenses to obscure or protect gradients. However, AttentionBreaker’s gradient-free capability bypasses these gradient-based defenses entirely by relying on magnitude-based importance, showing its potential to circumvent gradient-based restrictions. This adaptability highlights AttentionBreaker’s broad applicability in gradient-restricted environments, establishing it as a resilient approach for targeted model attacks.

### v3

@@ Abstract
Current BFA techniques fail to exploit this vulnerability due to the difficulty of traversing the large parameter space to identify critical parameters. To address this, we propose AttentionBreaker, a novel framework specifically designed for LLMs that facilitates efficient parameter-space traversal to identify critical parameters. Additionally, we introduce GenBFA, an evolutionary optimization strategy, to further narrow down the critical parameter set and identify the most vulnerable bits, facilitating a more efficient attack. Empirical findings indicate the pronounced vulnerability of LLMs to AttentionBreaker: for instance, perturbing merely $3$ bits ($4.129\times 10^{-9}\%$ of total parameters) in LLaMA3-8B 8-bit weight quantized (W8) model results in total performance collapse, with accuracy on Massive Multitask Language Understanding (MMLU) tasks plummeting from $67.3\%$ to $0\%$ and perplexity escalating from $12.6$ to $4.72\times 10^{5}$. These results highlight the framework’s efficacy in exploiting and exposing inherent vulnerabilities in LLM architectures. Code is open sourced at: https://github.com/TIES-Lab/attnbreaker.
@@ 6.1 Experimental Setup
Figure 6: Comparison of BFA attack using [12] and AttentionBreaker on LLaMA3-8B-Instruct W8, W4, and W1.58 models. Metrics are calculated on the “Astronomy” task of the MMLU benchmark.
@@ 6.2.1 AttentionBreaker on 8-bit Quantized LLMs
Here, we present the efficacy of the proposed AttentionBreaker framework on 8-bit weight quantized (W8) LLMs. Table II compares the effects of existing Bit-Flip Attack [12] and AttentionBreaker on 8-bit versions of LLM across several benchmarks. For each model, the precision, number of parameters, and bit-flip Count required to disrupt performance are listed in the Table. Benchmark results show that AttentionBreaker consistently reduces accuracy to zero across datasets, proving more efficient than existing BFA in degrading model performance with fewer bit-flips. The results highlight the impact of bit-flips on key model performance metrics, including perplexity, MMLU, and LM harness benchmarks.
For instance, LLaMA3-8B-Instruct W8 suffers accuracy drop to 00% on MMLU tasks from the original accuracy of $67.3$% by merely 3 bit-flips ($\mathbf{4.129\times 10^{-9}}$bit-flip rate) as captured in Figure LABEL:fig:LLaMA3-8Bwsweep. At the same time, the perplexity measure increases from $12.14$ to $4.72\times 10^{5}$ post-attack. In comparison, [12] require $\textbf{10}^{\textbf{7}}\times$ ($\approx\textbf{0.1}$% bit-flip rate) more bit-flips to accomplish a similar level of model degradation. A similar trend is observed for Phi-3-mini-128k-Instruct W8 models as well.
@@ 6.2.2 AttentionBreaker on 4-bit quantized LLMs
In this section, we present the performance of AttentionBreaker on 4-bit weight quantized (W4) LLMs. As shown in Figure LABEL:fig:llava_bfa, the attack on LLaMA3-8B W4 demonstrates that $\mathbf{7.8125\times 10^{-8}}$or 28 bit-flips results in a perplexity increase from $12.60$ to $3.8$×$\times 10^{4}$, with the MMLU score dropping from $60.3$% to 00. A similar trend is observed for Phi-3-mini-128k-Instruct W4 models as well. This showcases the efficiency of the proposed AttentionBreaker framework.
@@ 6.2.3 AttentionBreaker on low-precision LLM
We further evaluate the effectiveness of our approach for ultra-low precision models, such as BitNet W1.58[26]. BitNet W1.58 represents a 1.58-bit variant of LLMs where each model parameter (or weight) is constrained to ternary values $\{-1,0,1\}$. Given their low precision, these models are intuitively expected to be inherently resistant, as a single bit-flip induced error has limited impact. Employing our approach, we observe that 45811 or $45811\text{ or }5\times 10^{-4}$bit-flip rate is required to degrade the model performance to near 00% effectively as demonstrated in Figure LABEL:fig:bitnetwsweep. In comparison, [12] require $10^{5}\times$ ($\approx 1$% bit-flip rate) more bit-flips to accomplish a similar level of model degradation. This demonstrates the efficacy of AttentionBreaker in efficiently attacking low-precision models as well.
@@
Model Model type Precision Number of Parameters BFA[12] (# of bit-flips) AttentionBreaker (# of bit-flips) WikiText-2 MMLU LM Harness VQAv2 TextVQA
@@ 6.2.4 AttentionBreaker on VLM
We demonstrate the efficacy of AttentionBreaker on multimodal models such as VLMs, which integrate a vision encoder with a language model to enable multimodal tasks, including image comprehension, alongside text-based tasks. When applying our attack methodology to the LLaVA1.6-7B W8 model, which comprises 7.57 billion parameters, we observe that a minimal bit-flip rate of $\mathbf{2.394\times 10^{-8}}$%—equivalent to only 15 bit-flips—causes a substantial degradation in performance. Specifically, the VQAv2 and TextVQA scores, initially at $81.4\%$ and $64.2\%$, respectively, drop to $0\%$ following the attack. In comparison, [12] require ($\mathbf{10^{7}\approx 0.12}$bit-flip rate) more bit-flips to accomplish a similar level of model degradation. This result underscores the efficacy of the AttentionBreaker framework in exploiting latent vulnerabilities within multimodal LLMs as well.
@@ 6.4.1 Gradient-free AttentionBreaker
In this experiment, we investigate the effectiveness of AttentionBreaker in a gradient-free context, where gradient information is inaccessible at all stages. Here, the importance scores are computed solely based on weight magnitudes by setting $\alpha=0.0$. Our results demonstrate that this magnitude-based approach maintains the attack’s efficacy, as it accurately identifies critical weights, enabling a highly impactful attack. We observe that, without gradient information, setting $\alpha=0$ allows the model to be attacked with only $\mathbf{4.129\times 10^{-9}}\%$ bit-flips, reducing model accuracy to zero. This degradation matches the attack’s effectiveness at $\alpha=0.5$, as depicted in Figure LABEL:fig:LLaMA3-8Bwsweep. Subsequent analysis of the attack vector confirms that identical weights and bits are selected for attack across both settings, underscoring AttentionBreaker’s versatility and adaptability. This robustness is particularly significant as many defenses against BFAs assume that gradient access is essential for an attacker, prompting defenses to obscure or protect gradients. However, AttentionBreaker’s gradient-free capability bypasses these gradient-based defenses entirely by relying on magnitude-based importance, showing its potential to circumvent gradient-based restrictions. This adaptability highlights AttentionBreaker’s broad applicability in gradient-restricted environments, establishing it as a resilient approach for targeted model attacks.